# Is My Data in Your Retrieval Database? Membership Inference Attacks Against Retrieval Augmented Generation

### v2

@@ Abstract
Keywords:
AI Privacy; Membership Inference; RAG; Large Language Models
@@ 1 Introduction
Retrieval Augmented Generation (RAG) systems [12, 6] have emerged as a promising approach in natural language processing, gaining significant attention in recent years due to their ability to generate high-quality, up-to-date and contextually relevant responses. These systems combine retrieval and generation components to provide more accurate and informative responses compared to traditional language models. However, like any advanced technology, RAG systems are not immune to vulnerabilities.
@@
While previous research has successfully demonstrated various types of attacks against RAG systems [10, 21, 7, 20], there are still unexplored vulnerabilities in these systems that may pose privacy risks. Specifically, the use of the retrieval database introduces new privacy concerns. Since the retriever component searches a database for relevant passages, an attacker may be able to infer whether a certain text passage appears in the database by observing the outputs of the RAG system. This type of attack is known as a Membership Inference Attack (MIA), and can be used to reveal sensitive information about the contents of the retrieval database.
@@
MIAs were extensively researched in the past in the context of various machine and deep learning systems [15, 3, 2, 18, 9], however to the best of our knowledge the topic of membership inference against RAG systems remains under-explored.
@@
When a MIA is performed against a RAG system it can potentially reveal sensitive or proprietary company information. This may include information about individuals or organizations included in the retrieval database. Furthermore, MIA can be used to prove the unauthorized use of proprietary documents, as part of a legal action [16]. This dual capability makes them a particularly serious form of attack that must be investigated in order to ensure the security and privacy of these systems.
@@
In this study, we introduce a method that is both efficient and easy to use for conducting MIAs against RAG systems, with the objective of determining whether a specific data sample is present in the retrieval database. We assess the effectiveness of our attack using two benchmark datasets that represent scenarios where privacy may be a concern, and employing multiple generative models. Moreover, we introduce an initial defense strategy based on adding instructions to the RAG template to prevent the model from responding to the attack prompt. This approach seems to be highly effective for some datasets and models, but should be further developed to provide a more comprehensive solution.
@@ 2.1 Retrieval Augmented Generation
Retrieval Augmented Generation is a technique for enriching the knowledge of Large Language Models (LLMs) with external databases without requiring any additional training. This allows easy customization of trained LLMs for specific needs, such as creating AI-based personal assistants or making long textual content (such as a user manual) accessible using simple text queries [4].
Prior to deploying a RAG pipeline, a retrieval database $\mathcal{D}$ must be populated with documents. During this initialization phase, each document is split into chunks, which are mapped into vector representations using a retrieval model $E$, and then stored as an index in the database alongside the original document.
The retrieval model $E$ is specifically designed to learn a mapping from user prompts and documents to a shared vector space, such that prompts are embedded close together with documents containing relevant information to respond to the prompt. This enables efficient searching in the retrieval database, as semantically similar prompts and documents are clustered together in the vector space.
Once deployed, the RAG pipeline typically consists of two main phases: search and generation. In the search phase, $\mathcal{D}$ is queried to find relevant documents that match the user’s query or prompt. When a user prompt is processed by the RAG pipeline, the prompt $p$ is mapped into a vector representation using $E$. Then, $\mathcal{D}$ is searched to find the top-k most similar entries, based on a distance metric calculated using the vector representations (e.g., Euclidean distance). The retrieved entries from $\mathcal{D}$ are organized and provided to the generation phase together with the user prompt.
@@
In practice, it is common to place titles indicating the different areas in the input or additional instructions. For example, placing the sentence ”Answer the question based on the context” after the context and before the user prompt. A full example of a RAG template appears in Figure 1.
@@ 2.2 Membership Inference Attacks
Membership inference attacks [15, 9] are a type of privacy threat, where an attacker aims to determine whether a specific data record was used in the training set of a machine learning model. This carries significant privacy implications as it can potentially reveal sensitive information about individuals, even if the model does not directly release any personal data.
Formally, an attacker aims to determine the membership of a sample $x$ in the training data of a target model $\mathcal{D}_{m}$, i.e., to check if $x\in\mathcal{D}_{m}$, which is known as sample-level membership inference. Typically, these attacks involve calculating one or more metrics on the target model’s outputs that reflect the probability of the sample being a part of the training set, such as the model outputs’ entropy or log-probabilities [3]. Several metrics may be computed for each sample and then fused together using a machine learning model, known as an attack model, which in turn outputs the probability of a sample being a member of the training set. Additionally, an attacker may also aim to determine the membership of a certain user, i.e., to check if a user’s data is part of the training set, which is known as user-level membership inference [14]. Throughout this paper we will address the membership inference challenge from a sample-level perspective.
@@ 2.3 Threat Model
In addition to the black-box setting, we also evaluate a supplementary gray-box scenario, in which the attacker has access to the log-probabilities of the generated tokens. Moreover, in this scenario, we assume that the attacker can train the attack model on the target model’s outputs on a subset of the model’s actual training and test datasets [13].
@@ 3 Methodology
Figure 2: Attack prompt example for RAG MIA. The highlighted text is the attack-specific part of the prompt, and the rest is taken from the sample for which membership is inferred.
In this section we present our membership inference attack. The input to the generative model used by the system consists of the RAG tempate into which the retrieved database entries and the user prompt are inserted, as depicted in Figure 1. Since the attacker’s interaction with the system is limited to the user prompt part of the RAG template, the attack must be based on a prompt that achieves both of the following goals:
@@
The first goal can be easily fulfilled by creating a prompt that has high similarity with the desired sample in the retrieval database. Essentially, any prompt that contains the target sample should be sufficient, as long as there is not too much text added to it. However, the prompt must also cause the generation process to expose the membership status of the sample. Prompting the system with target sample alone will not necessarily achieve this desired result.
To handle this, we designed the following attack prompt format, along with a few additional variations presented in Table 1: ”Does this: ”{Target Sample}” appear in the context? Answer with Yes or No.”. In Figure 2 we present an example of this attack prompt format. Using this prompt we are able to achieve both goals: cause the right sample to be retrieved from the retrieval database, and make the generation phase produce a binary response regarding the sample’s presence in the context. In the black-box setting, we use the model’s answer (Yes/No) alone to deduce the membership status of samples.
As an enhancement to our attack, in cases where the adversary has access to the log-probabilities of the selected tokens, we additionally employ an attack model, similar to the approach described in Section 2.2. In this setup, we apply an ensemble of attack models [13], using as input to the attack both the logits and the class-scaled-logits [3] corresponding to the ”Yes” and ”No” tokens output from the attacked model. The logits are computed by first calculating the exponent of of the log-probability to get a probability estimate $P$ and then applying the logit function.
Since the model only outputs the log-probability of the selected token, for example the ”Yes” token, without the complementary ”No” token, we assign a fixed low probability value of $0.001$ to the complementary token.
Black-Box
Gray-Box
Figure 3: Comparison of different attack prompts. Top: HealthCareMagic, Bottom: Enron
@@ 3.1 Experimental Setup
  • google/flan-ul2 [17], denoted by flan
  • meta-llama/llama-3-8b-instruct [1], denoted by llama
  • mistralai/mistral-7b-instruct-v0-2 [11], denoted by mistral
@@
We used a Milvus Lite [19, 8] vector database with $k=4$, Euclidean distance (L2) metric type and HNSW index.
@@
Please answer the question using the context provided.
If the question is unanswerable, say unanswerable”.
@@
In our evaluation we experimented with 5 different attack prompts, listed in Table 1. Each attack prompt includes a placeholder for a sample, which can be a member or a non-member sample, as shown in Section 3. In the case of the Enron dataset, the sample is the full email body, or its first 1000 characters if it is longer. In HealthCareMagic, the human part of the dialogue is used as the sample.
@@
In both the black-box and the gray-box scenarios we ran our attack using random samples containing 2000 member documents and 2000 non-member documents. In the gray-box case, these documents were split between 40 attack models in the ensemble [13], each model trained 5 times with different train-test splits. In this evaluation, we classified model outputs that failed to include either the ”Yes” or ”No” tokens as non-member documents. This case accounted for approximately 6% of the total outputs. We present the statistics on the amount of such documents in Appendix C.
@@ 4.1 Impact of Attack Prompts
One important aspect of performing attacks against any generative model is the prompt used to trigger the attack. In the case of RAG, the attack prompt is first used to fetch the relevant documents from the retrieval database, and is then incorporated together with the fetched documents into the RAG template and provided to the generative model. To achieve the best performance, we experimented with five possible attack prompts, described in Table 1. Each one tries to manipulate the RAG system in a slightly different manner to both fetch the relevant documents from the retrieval database, and cause the generative model to output whether the target document was indeed in the retrieval database.
@@
Unsurprisingly, the results in the black-box setting are inferior to those in the gray-box setting. This is especially the case for the flan model, with an improvement of up to 25% in the gray-box setting. This means that the log-likelihood values for member documents are significantly higher than for non-member documents, indicating a higher confidence of the model in its response. However, when looking at the llama and mistral models, the average difference is only up to 12%, and in some cases even lower, depending on the prompt.
To further explore this difference, we analyze the percentage of samples that are correctly retrieved from the database for each prompt. We found that over 95% of the member samples are indeed retrieved for both datasets. This is in contrast with the non-member samples, which are retrieved in nearly 0% of the cases. The full results of this analysis can be found in Appendix B. Thus, we conclude that the flan model is more grounded to the content of its input prompt (context grounding), and thereby more sure of the presence/absence of a piece of text from it in comparison to the llama and mistral models.
@@ 4.2 Attack Results Summary
To quantify the overall MIA risk, we summarize the best results (across prompts) for each combination of model and dataset in Table 2 . We can see that, on average, the attack success rate is highest for flan, in both threat models. In addition, we observe that the overall risk in the black-box setting is similar between almost all models, with the exception of the HealthCareMagic-llama case. However in the gray-box setting the results vary more.
Table 2: RAG-MIA results summary
@@
HealthCareMagic flan 0.80 1.00
llama 0.88 0.96
mistral 0.74 0.83
Enron flan 0.82 0.96
llama 0.79 0.83
mistral 0.78 0.82
Our lowest attack AUC-ROC score of 0.74 is quite high compared to previous research on sample-level MIA in language models [5, 14]. Overall, our black-box attack achieves an average AUC-ROC of 0.80 and the gray-box attack achieves an average AUC-ROC of 0.90. The best gray-box attack on flan accomplishes almost perfect performance. These results highlight the risk associated with deploying RAG based systems without proper defense mechanisms.
4.3 Defense Strategy
To counter the threat of RAG-MIA attacks, we propose an initial defense strategy in this section. Building on previous research that has demonstrated the ability to influence the output of LLMs through system instructions, we leverage this concept to develop a defense mechanism. Our approach involves modifying the RAG template to include explicit instructions for the LLM to disregard requests that directly inquire about the retrieved database content, thereby providing a defense against RAG-MIA attacks.

### v3

@@ 1 INTRODUCTION
Retrieval Augmented Generation (RAG) systems [Lewis et al., 2020, Gao et al., 2023] have emerged as a promising approach in natural language processing, gaining significant attention in recent years due to their ability to generate high-quality, up-to-date and contextually relevant responses. These systems combine a retrieval database and retrieval and generation components to provide more accurate and informative responses compared to traditional language models. For example, in the health domain, access to timely medical literature, research papers and patient records using a RAG architecture can assist in expediting diagnosis and improving patient outcomes. However, like any advanced technology, RAG systems are not immune to vulnerabilities.
@@
While previous research has successfully demonstrated various types of attacks against RAG systems [Hu et al., 2024, Zou et al., 2024, Greshake et al., 2023, Zeng et al., 2024], there are still unexplored vulnerabilities in these systems that may pose privacy risks. Specifically, the use of a retrieval database, that may contain sensitive or proprietary information, introduces new privacy concerns for the data residing in that database. Since the retrieval database is searched for relevant passages to aid the model in responding to a specific user prompt, an attacker may be able to infer whether a certain text passage appears in the database by observing the outputs of the RAG system. This type of attack is known as a Membership Inference Attack (MIA), and can be used to reveal sensitive information about the contents of the retrieval database.
@@
MIAs were extensively researched in the past in the context of various machine and deep learning models [Shokri et al., 2017, Carlini et al., 2022, Amit et al., 2024, Tseng et al., 2021, Hu et al., 2022], facilitating the detection of models’ training data. However, to the best of our knowledge, the topic of membership inference against RAG systems remains under-explored.
@@
When a MIA is performed against a RAG system, it can potentially reveal sensitive or proprietary company information, including information about individuals or organizations included in the retrieval database. Furthermore, MIA can be used to prove the unauthorized use of proprietary documents, as part of a legal action [Song and Shmatikov, 2019]. This dual capability makes them a serious form of attack that must be investigated to ensure the security and privacy of these systems. However, existing approaches to performing MIA on LLMs may not necessarily succeed in leaking membership information about RAG documents. Specifically, we show that careful selection of the attack prompt, and adapting it to the RAG scenario, is crucial to its success.
In this study, we introduce a method that is both efficient and easy to use for conducting MIAs against RAG systems, with the objective of determining whether a specific data sample is present in the retrieval database. The method treats the entire RAG pipeline as a black box and does not require access to, or knowledge of, any internal implementation details, or even the RAG template, making it especially useful and easy to employ.
We assess the effectiveness of our attack using two benchmark datasets that represent various domains where privacy may be a concern, namely medical questions and answers and emails, and employing multiple generative models.
@@
Moreover, we introduce an initial defense strategy based on adding instructions to the RAG template to prevent the model from responding to the attack prompt. This approach does not require introducing any additional components or add any computational overhead to the overall RAG solution. This defense seems to be highly effective for some datasets and models, but should be further developed to provide a more comprehensive solution.
@@ 2.1 Retrieval Augmented Generation
Retrieval Augmented Generation is a technique for enriching the knowledge of Large Language Models (LLMs) with external databases without requiring any additional training. This allows easy customization of trained LLMs for specific needs, such as creating AI-based personal assistants or making long textual content (such as a user manual) accessible using simple text queries [Chase, 2022].
Prior to deploying a RAG pipeline, a retrieval database $\mathcal{D}$ must be populated with documents. During this initialization phase, each document is split into chunks, which are mapped into vector representations (embeddings) using an embedding model $E$, and then stored as an index in the database alongside the original document.
The embedding model $E$ is specifically designed to learn a mapping from user prompts and documents to a shared vector space, such that prompts are embedded close together with documents containing relevant information to respond to the prompt. This enables efficient searching in the retrieval database, as semantically similar prompts and documents are clustered together in the vector space.
Once deployed, the RAG pipeline typically consists of two main phases: search and generation. In the search phase, $\mathcal{D}$ is queried to find relevant documents that match the user’s query or prompt. When a user prompt is processed by the RAG pipeline, the prompt $p$ is mapped into a vector representation using $E$. Then, $\mathcal{D}$ is searched to find the top-k most similar entries, based on a distance metric calculated on the vector representations (e.g., Euclidean distance). The retrieved entries from $\mathcal{D}$ are organized and provided to the generation phase together with the user prompt.
@@
In practice, it is common to place titles in the template to indicate the different areas, as well as provide additional instructions. For example, placing the sentence ”Answer the question based on the context” after the context and before the user prompt. A full example of a RAG template appears in Figure 1.
@@ 2.2 Membership Inference Attacks
Membership inference attacks [Shokri et al., 2017, Hu et al., 2022] are a type of privacy threat, where an attacker aims to determine whether a specific data record was used in the training set of a machine learning model. This carries significant privacy implications as it can potentially reveal sensitive information about individuals, even if the model does not directly release any personal data.
Formally, an attacker aims to determine the membership of a sample $x$ in the training data $\mathcal{D}_{m}$ of a target model $m$, i.e., to check if $x\in\mathcal{D}_{m}$. This is known as sample-level membership inference. Typically, these attacks involve calculating one or more metrics on the target model’s outputs that reflect the probability of the sample being a part of the training set, such as the model outputs’ entropy or log-probabilities [Carlini et al., 2022]. Several metrics may be computed for each sample and then fused together using a machine learning model, known as an attack model, which in turn outputs the probability of a sample being a member of the training set.
Additionally, an attacker may also aim to determine the membership of a certain user, i.e., to check if a user’s data is part of the training set, which is known as user-level membership inference [Shejwalkar et al., 2021]. Throughout this paper we will address the membership inference challenge from a sample-level perspective.
@@ 2.3 Threat Model
In addition to the black-box setting, we also evaluate a supplementary gray-box scenario, in which the attacker has access to the log-probabilities of the generated tokens, as previously explored in prior art [Duan et al., 2024, Zhang et al., 2024] Moreover, in this scenario, we assume that the attacker can train the attack model on a subset of the model’s actual training and test datasets [Shachor et al., 2023].
3 RELATED WORK
3.1 Membership Inference Attacks against LLMs
Membership inference attacks have been extensively explored for various types of machine and deep learning models [Shokri et al., 2017, Carlini et al., 2022, Hu et al., 2022]. Recent interest in Large Language Models (LLMs), has brought a variety of MIA studies tailored for such models [Mahloujifar et al., 2021, Kandpal et al., 2023, Panda et al., 2024]. While some studies utilized existing concepts such as loss-based attacks [Shejwalkar et al., 2021], others employed domain-specific approaches, for example the use of the Perplexity measure to indicate membership [Galli et al., 2024].
After RAG was introduced to enhance the capabilities of LLMs, it also became a target for privacy attacks, such as [Zeng et al., 2024], that attempts to extract data from the retrieval database. However, to the best of our knowledge, this is the first paper that suggests a MIA against such systems. Shortly after our preliminary publication, another paper suggested a different approach to MIA against RAG systems, utilizing the semantic similarity between the model’s generated text and the original sample as the membership metric [Li et al., 2024b]. However, our method does not rely on the model to actually leak the original text, which is usually considered a harder problem than that of determining membership alone.
3.2 Prompt Injection Attacks
A prompt injection attack aims to compromise an LLM-based application so that it produces some desired response, for example, changing the LLM prediction or utilizing the LLM for another purpose that is not its intended task. Our attack can be considered a kind of prompt injection attack  [Liu et al., 2024], whose goal is to violate the privacy of the retrieval dataset.
Prompt injection attacks can take several forms. According to the framework described in  [Liu et al., 2024], our attack falls under the category of ”Naive attacks”, such as [Willison, 2022, Matthew Kosinski, 2024], in which the adversary’s instruction is simply concatenated to the target data.
Other types of prompt injection attacks have also been employed against RAG systems, with the goal of interfering with their outputs. In [Zhong et al., 2023], the authors describe a poisoning attack where the adversary introduces malicious documents into the retrieval database, causing the RAG system to output undesirable responses in response to certain queries.
@@ 4 METHODOLOGY
Figure 2: Overall Flow of our MIA Attack on a RAG pipeline.
Figure 3: Attack prompt example for RAG-MIA. The highlighted text is the attack-specific part of the prompt, and the rest is taken from the sample for which membership is inferred.
In this section we present our membership inference attack (denoted as RAG-MIA), which is shown in Figure 2. The input to the generative model consists of the RAG template into which the retrieved database entries and the user prompt are inserted, as depicted in Figure 1. Since the attacker’s interaction with the system is limited to the user prompt part of the RAG template, the attack must be based on a prompt that achieves both of the following goals:
@@
The first goal can be easily fulfilled by creating a prompt that has high similarity with the desired sample in the retrieval database. Essentially, any prompt that contains the target sample without too much additional text should be sufficient. However, the prompt must also cause the generation process to expose the membership status of the sample. Prompting the system with the target sample alone will not necessarily achieve this goal.
To handle this, we designed the following attack prompt format (along with a few additional variations presented in Table 1): ”Does this: ”{Target Sample}” appear in the context? Answer with Yes or No.”. In Figure 3 we present an example of this attack prompt format with a specific target sample. Using this prompt we are able to achieve both goals: cause the right sample to be retrieved from the retrieval database, and make the generation phase produce a binary response regarding the sample’s presence in the context. In the black-box setting, we use the model’s answer (Yes/No) alone to deduce the membership status of samples.
As an enhancement to our attack, in cases where the adversary has access to the log-probabilities of the selected tokens [Zhang et al., 2024, Duan et al., 2024], we additionally employ an attack model (see Section 2.2) to determine membership. In this setup, which we call the gray-box setting, we employ an ensemble of attack models [Shachor et al., 2023] that receive as input both the logits and class-scaled logits [Carlini et al., 2022] corresponding to the ”Yes” or ”No” token output from the target model. The logits are computed by first calculating the exponent of the log-probability to get a probability estimate $P$ and then applying the logit function.
Since the model only outputs the log-probability of the selected token, for example the ”Yes” token, without the complementary ”No” token, we assign a fixed low probability value of $0.001$ to the complementary token. More details about this attack setup can be found in Section MIA Attack Details.
@@ 4.1 Experimental Setup
We used a Milvus Lite [Wang et al., 2021, Guo et al., 2022] vector database with $k=4$, Euclidean distance (L2) metric type and HNSW index.
@@
Please answer the question using the context
provided. If the question is unanswerable, say
"unanswerable".
@@
In our evaluation we experimented with 5 different attack prompts, listed in Table 1. Each attack prompt includes a placeholder for a sample, which can be a member or a non-member sample. In the case of the Enron dataset, the sample is the full email body, or its first 1000 characters if it is longer. In the HealthCareMagic dataset, the human part of the dialogue is used as the sample.
@@
In both the black-box and the gray-box scenarios, we started by randomly sampling 2000 member documents and 2000 non-member documents to be used in the evaluation. We then ran our attack 10 times, each time using a different random sample of 500 member documents and 500 non-member documents out of those 2000 member or non-member documents selected in the previous stage. For the gray-box scenario we leveraged the attack technique proposed by [Shachor et al., 2023], which combines multiple attack models into a powerful ensemble. Each model in the ensemble is trained on a small subset of the data and optimized using a comprehensive hyper-parameter search covering various attack model types, data scaling approaches, and model training parameters. This creates diversity among model predictions, and improves the overall membership inference attack performance. In our experiments, we used an ensemble consisting of 40 attack models.
Throughout the experiments in this paper, model outputs that failed to include either ”Yes” or ”No” tokens were classified as non-member documents. This case accounted for approximately 6% of the total outputs. We present detailed statistics on this in Appendix C.
Evaluation Metrics
We employ three primary evaluation metrics to assess the performance of our attack. Firstly, for the gray-box attack, we employ TPR@lowFPR [Carlini et al., 2022], which measures the percentage of correctly identified member samples (True Positive Rate) at a fixed, low percentage of falsely identified member samples (False Positive Rate). In our case, we present results for the extreme case of FPR=0, where all non-member samples are correctly classified. Secondly, for the black-box attack, where we simply use the Yes/No answer generated by the model and have no threshold that can be tuned, we report the single TPR and FPR values achieved by the attack (denoted as TPR@FPR). Thirdly, we utilize the Area Under the Receiver Operating Characteristic curve (AUC-ROC), a common metric for evaluating binary classification problems, including membership inference attacks.
@@ 5 RESULTS
Enron Enron HealthcareMagic HealthcareMagic
AUC-ROC TPR@FPR AUC-ROC TPR@FPR
Black-Box
Refer to caption Refer to caption Refer to caption Refer to caption
Gray-Box
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Comparison of different attack prompts. The top row shows results for black-box attacks, evaluated using AUC-ROC and TPR@FPR. The wide bars show the TPR and the narrow bars inside them show the respective FPR values. The bottom row shows results for gray-box attacks, evaluated using AUC-ROC and TPR@lowFPR.
@@ 5.1 Impact of Attack Prompts
One important aspect of performing attacks against any generative model is the prompt used to trigger the attack. In the case of RAG, the attack prompt is first used to fetch the relevant documents from the retrieval database, then it is incorporated into the RAG template together with the fetched documents and used as input for the generative model. To achieve the best performance, we experimented with five possible attack prompts, detailed in Table 1. Each one tries to manipulate the RAG system in a slightly different manner to both fetch the relevant documents from the retrieval database, and cause the generative model to output whether the target document was indeed in the retrieval database.
@@
Answer with one word. Answer No if you are not sure.
In Figure 4 we present the AUC-ROC and TPR results for both the gray-box and black-box scenarios, measuring the effectiveness of the different attack prompts. In the black-box scenario, the attack predictions are discrete, meaning that computing TPR@lowFPR is not possible. Instead we present the TPR@FPR of the attack, i.e., the percentage of members correctly classified as members and the percentage of non-members classified as members. The same results organized in table format can be found in Appendix 7.
@@
Unsurprisingly, the TPR@FPR and the AUC-ROC results in the gray-box setting are superior to those in the black-box setting. This is prominent in the case of the flan model, with an improvement of up to 22% in the gray-box setting. This means that for this model, the log-probability values for member documents are significantly higher than for non-member documents, indicating a higher confidence of the model in its response. However, when looking at the llama and mistral models, the average difference is only up to 7%, and in some cases even lower, depending on the prompt.
To further explore this difference between the models, we analyzed the percentage of member samples that are correctly retrieved from the database for each prompt. We found that over 95% of the member samples are indeed retrieved for both datasets. This is in contrast with the non-member samples, that are retrieved in nearly 0% of the cases. The full results of this analysis can be found in Appendix B. Thus, we conclude that the flan model is more grounded to the content of its input prompt (i.e., context grounded), and thereby more sure of the presence/absence of a piece of text from it in comparison to the llama and mistral models.
@@ 5.2 Attack Results Summary
Table 2: RAG-MIA results summary.
@@
HealthCareMagic flan 1.00 0.61 0.85 0.81 0.99
llama 0.95 0.20 0.73 0.89 0.96
mistral 0.42 0.10 0.36 0.74 0.83
Enron flan 1.00 0.56 0.63 0.82 0.96
llama 0.78 0.30 0.28 0.79 0.83
mistral 0.61 0.17 0.22 0.78 0.81
We present a summary of the best results from Figure 4 for each model/dataset combination in Table 2, which quantifies the overall risk of Membership Inference Attacks.
The results show that the attack is most effective against the flan models, in both the gray- and black-box scenarios. Notably, the overall risk in both scenarios is very high, reaching nearly perfect AUC-ROC in the gray-box scenario (with an average of 0.9 across models and datasets) and close to 0.9 AUC-ROC in the black-box scenario (0.8 on average). The TPR@lowFPR values are also remarkably high, ranging from 0.22 to 0.85 (0.51 on average) for an FPR of 0. This significantly surpasses TPR results from previous MIA research in language models [Carlini et al., 2022, Li et al., 2024a, Zhang et al., 2024], which is usually no higher than 0.25 even for an FPR of 0.05. In the black-box scenario, the TPR for all cases is higher by at least 30% than the FPR. This suggests that the attack is highly effective in real-world scenarios where the attacker aims to correctly identify member samples.
These results underscore the significant risk associated with deploying RAG-based systems without adequate defense mechanisms in place.
5.3 Initial Defense Strategy
To counter the threat of RAG-MIA attacks, in this section we propose an initial defense strategy. Building on previous research that has demonstrated the ability to influence the output of LLMs through system instructions [Reynolds and McDonell, 2021, Lyu et al., 2024], we leverage this concept to develop a basic defense mechanism. Our approach involves modifying the RAG template to include explicit instructions for the LLM to disregard requests that directly inquire about the retrieved database content.